Question: A company currently pays a dividend of $1.25 per share (D0 = $1.25). It is estimated that the company's dividend will grow at a rate of 16% per year for the next 2 years, and then at a constant rate of 8% thereafter. The company's stock has a beta of 1.95, the risk-free rate is 4%, and the market risk premium is 3%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent.

Tried to figure it out with other solutions, yet it still didn't make sense. Thanks in advance!
Answer #1

Required Return, rs = Risk-free Rate + Beta * Market Risk Premium
Required Return, rs = 4.00% + 1.95 * 3.00%
Required Return, rs = 9.85%

Last Dividend, D0 = $1.25

Growth rate for next 2 years is 16%, and a constant growth rate (g) of 8% thereafter

D1 = $1.2500 * 1.16 = $1.4500
D2 = $1.4500 * 1.16 = $1.6820
D3 = $1.6820 * 1.08 = $1.81656

P2 = D3 / (rs - g)
P2 = $1.81656 / (0.0985 - 0.08)
P2 = $98.1924

P0 = $1.45/1.0985 + $1.682/1.0985^2 + $98.1924/1.0985^2
P0 = $84.09

Current stock price is $84.09